# Phase space and Data analyses of a non-minimally coupled scalar field system with decaying dark energy model

### v1

@@ VII Conclusion
In this paper, we used Q-SC-CDM model to study the cosmological dynamics with non-minimal coupling. We started the detailed investigations of the dynamics with a well known example of $F(\phi)R$ coupling, where $F(\phi)=1-\xi\phi^{2}$. The numerical evolution of Q-SC-CDM potential was shown in Fig. 1. Initially, the potential $V(\phi)$ exhibited the positive behavior for negative values of scalar field. As field evolved, the universe obtained an accelerated expansion and remained so untill $\phi_{dec}$. The potential showed negative behavior on the further evolution of field that lead to the collapse of the universe, referred to as Big Crunch Singularity. We numerically checked the range of $\xi$ to avoid such singularity. It was noticed that the Big Crunch Singularity remained present in the range of $0\leq\xi\leq 0.3$ whereas it was completely vanished for $\xi>0.3$. In our work, we chose two distinct values of $\xi$ such as 0.1 and 0.5. For $\xi\rightarrow 0$, NMC scalar field model was reduced to the standard quintessence and Big Crunch Singularity remained present. In case of $\xi=0.1$, we numerically evolved the evolution equations (14) and (15). The results were depicted in Fig. 2. The numerical evolution of the universe is divided into three distinct regions: decelerated expansion, accelerated expansion and slow-contraction. In this case, Big Crunch Singularity is delayed in distant future but not completely avoided, see Fig. 2. For $\xi=0.5$, the slow-contraction phase vanished. Consequently, the Big Crunch Singularity was redundant. Therefore, we obtained an accelerated expansion for a longer period and de-Sitter expansion in distant future. The numerical evolution of different cosmological parameters were shown in Fig. 3 which were well approximated by $\Lambda$CDM model.
@@
Next, we constructed an autonomous system by using a convenient set of dimensionless variables to performed the phase space analysis. We found stationary points, discussed their stabilities by looking at the eigenvalues, and asymptotic regimes of solutions in the model. The critical point $A_{1}$ behaved as a saddle, and corresponded to a non-accelerating universe with $\Omega_{\phi}=0$ and $w_{eff}=0$. As a result, the scalar field was not contributing to the energy density at this point. The fixed point $A_{2}$ was also exhibited saddle behavior with two eigenvalues positive and the other two negative. The fixed points $A_{3}$ and $A_{4}$ were behaved as a stable node, and exhibited accelerated expansion with $\Omega_{\phi}=1$ and $w_{eff}=w_{\phi}=-1$, which mimicked the cosmological constant-like behavior. The phase portraits for both points were shown in Fig. 4, where all trajectories approached the stable attractor points.
@@
Finally, we constrained the model parameters by fixing the non-minimal coupling parameter $\xi$ to two representative values: $\xi=0.1$ and 0.05. These values were chosen to avoid numerical stiffness encountered while solving the equations. We also observed that increasing $\xi$ beyond 0.1 leads to divergence in other dynamical variable, making the numerical evolution extremely difficult. Therefore, our comparison is restricted to these two cases. The models were tested using two distinct supernova datasets—Pantheon+ (PP) and the Dark Energy Survey (DES)—while keeping the base datasets (CC + DESBAO) fixed. The results indicate that both models exhibit similar behavior when applied to the same dataset and show a notable departure from the standard $\Lambda$CDM model at higher redshifts. This deviation is characterized using the $Om(z)$ diagnostic, which reveals that for the PP dataset, the effective fluid energy density in both models resembles that of a cold dark matter component. However, for the DES dataset, the deviation becomes significantly more pronounced.
Additionally, we assessed the statistical significance of the models relative to $\Lambda$CDM using the AIC and BIC. The AIC results indicate moderate support for both models, whereas the BIC suggests only weak support in comparison to the fiducial model.
@@ Acknowledgments
This research was funded by the Science Committee of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP22682760).
SH acknowledges the support of National Natural Science Foundation of China under Grants No. W2433018 and No. 11675143, and the National Key Research and development Program of China under Grant No. 2020YFC2201503.
MS acknowledges Integral University, Lucknow for financial support through Seed Money Grant 2024-2025 (Project Sanction No.: IUL/ICEIR/SMP/2024-04) and MCN: IU/R&D/2025-MCN000…. MS also thanks the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune for the hospitality and facilities under the visiting associateship program where the work was completed.

### v2

@@ VII Conclusion
In this paper, we used Q-SC-CDM model to study the cosmological dynamics with non-minimal coupling. We started the detailed investigations of the dynamics with a well known example of $F(\phi)R$ coupling, where $F(\phi)=1-\xi\phi^{2}$. The numerical evolution of Q-SC-CDM potential was shown in Fig. 1. Initially, the potential $V(\phi)$ exhibited the positive behavior for negative values of scalar field. As field evolved, the universe obtained an accelerated expansion and remained so untill $\phi_{dec}$. The potential showed negative behavior on the further evolution of field that led to the collapse of the universe, referred to as Big Crunch Singularity. We numerically checked the range of $\xi$ to avoid such singularity. It was noticed that the Big Crunch Singularity remained present in the range of $0.20\leq\xi<$ 0.2 whereas it was completely vanished for $\xi\geq 0.2$. In our work, we chose two distinct values of $\xi$ such as 0.1 and 0.5. For $\xi\rightarrow 0$, NMC scalar field model was reduced to the standard quintessence and Big Crunch Singularity remained present. In case of $\xi=0.1$, we numerically evolved the evolution equations (14) and (15). The results were depicted in Fig. 2. The numerical evolution of the universe was divided into three distinct regions: decelerated expansion, accelerated expansion and slow-contraction. In this case, Big Crunch Singularity is delayed in distant future but not completely avoided, see Fig. 2. For $\xi=0.5$, the slow-contraction phase vanished. Consequently, the Big Crunch Singularity was redundant. Therefore, we obtained an accelerated expansion for a longer period and de-Sitter expansion in distant future. The numerical evolution of different cosmological parameters were shown in Fig. 3 which were well approximated by $\Lambda$CDM model.
@@
Next, we constructed an autonomous system by using a convenient set of dimensionless variables to performed the phase space analysis. We obtained stationary points, discussed their stabilities by looking at the eigenvalues, and found asymptotic regimes of solutions in the model. The critical point $A_{1}$ behaved as a saddle, and corresponded to a non-accelerating universe with $\Omega_{\phi}=0$ and $w_{eff}=0$. As a result, the scalar field was not contributing to the energy density at this point. The fixed point $A_{2}$ was also exhibited saddle behavior with two eigenvalues positive and the other two negative. The fixed points $A_{3}$ and $A_{4}$ were behaved as a stable node, and exhibited accelerated expansion with $\Omega_{\phi}=1$ and $w_{eff}=w_{\phi}=-1$, which mimicked the cosmological constant-like behavior. The phase portraits for both points were shown in Fig. 4, where all trajectories approached the stable attractor points.
@@
Finally, we constrained the model parameters by fixing the non-minimal coupling parameter $\xi$ to two representative values: $\xi=0.1$ and 0.05. These values were chosen to avoid numerical stiffness encountered while solving the equations. We also observed that increasing $\xi$ beyond 0.1 leads to divergence in other dynamical variable, making the numerical evolution extremely difficult. Therefore, our comparison is restricted to these two cases. The models were tested using two distinct supernova datasets—Pantheon+ and DES—while keeping the base datasets (CC + DESBAO) fixed. The results indicate that both models exhibit similar behavior when applied to the same dataset and show a notable departure from the standard $\Lambda$CDM model at higher redshifts. This deviation is characterized using the $Om(z)$ diagnostic, which reveals that for the PP dataset, the effective fluid energy density in both models resembles that of a cold dark matter component. However, for the DES dataset, the deviation becomes significantly more pronounced. Additionally, we assessed the statistical significance of the models relative to $\Lambda$CDM using the AIC and BIC. The AIC results indicate moderate support for both models, whereas the BIC suggests only weak support in comparison to the fiducial model.
@@ Acknowledgments
This research was funded by the Science Committee of the Ministry of Science and Higher Education of the Republic of Kazakhstan (Grant No. AP22682760). SH acknowledges the support of National Natural Science Foundation of China under Grants No. W2433018 and No. 11675143, and the National Key Research and development Program of China under Grant No. 2020YFC2201503. MS acknowledges Integral University, Lucknow for financial support through Seed Money Grant 2024-2025 (Project Sanction No.: IUL/ICEIR/SMP/2024-04) and MCN: IU/R&D/2025-MCN0003753. MS also thanks the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune for the hospitality and facilities under the visiting associateship program where the work was completed.